OpenStudy (anonymous):

A man earns $112 interest in one year on an investment of $1,400. What was his rate of interest?

OpenStudy (ranga):

Interest = P * r * t = $112 where P is the principal = $1,400 r is the rate of interest per year in decimal. t is the number of years invested = 1 year Substitute the numbers in (1) and solve for r. Multiply by 100 to convert decimal to percentage.

OpenStudy (ranga):

The first line is equation (1)

OpenStudy (anonymous):

look i dont even know how to multiply very good let alone all of this...

OpenStudy (ranga):

Interest = P * r * t = 112 Put P = 1400, t = 1 in the above equation: 1400 * r * 1 = 112 divide both sides by 1400 r = 112/1400 = 0.08 Multiply 0.08 by 100 to convert it to percentage. r = 0.08 * 100 = 8%